Doc. Math. J. DMV 631

Appendix: Invited Forty-Five Minute Lectures at the Section Meetings

This appendix contains three manuscripts of Invited Speakers which are not included in Volume II or III

Section 1. Logic:

A. J. Wilkie: O-Minimality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I 633

Section 10. Partial Differential Equations:

Mikhail Safonov: Estimates Near the Boundary for Solutions of Second Order Parabolic Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I 637

Section 12. Probability and Statistics:

F. Götze: Errata: Lattice Point Problems . . . . . . . . . . . . . . . . . . . . . . . . I 648

Section 16. Applications:

Serafim Batzoglou, Bonnie Berger, Daniel J. Kleitman, Eric S. Lander, and Lior Pachter: Recent Developments in Computational Gene Recognition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I 649

Documenta Mathematica · Extra Volume ICM 1998 · I